# Hacking Cryptographic Protocols with Tensor Network Attacks

### v1

@@ I Introduction
Moreover, in the NISQ era, simulations can offer reliable benchmarks of noisy quantum computers Patra et al. (2024a), which motivates us further to use our quantum-inspired tensor network methods. Such methods have also seen substantial advancements Orús (2014). Tensor networks such as Matrix Product States (MPS) have proven effective in simulating quantum systems Orús (2019); Verstraete et al. (2008). MPS are especially powerful in modeling one-dimensional systems, efficiently capturing their entanglement structure with polynomial resources. This efficiency stands in contrast to the exponential resources required by brute-force methods. The success of MPS in one-dimensional systems has inspired extensions to higher dimensions, such as Projected Entangled Pair States (PEPS), which handle two-dimensional systems and beyond. As an example, we recently introduced the so-called Flexible-PEPS approach Patra et al. (2024b), able to simulate quantum complex systems without a geometry constraint. When simulating quantum computers, the resulting Flexible-PEPS Quantum Circuit Simulator (FQCS) is particularly powerful.
@@ II.1 Variational Quantum Attack
where the columns of matrix $A$, i.e., $a_{i}$, represent the $k$ maximally orthogonal states over $N$ qubits and $\|\cdot\|_{F}$ denotes the Frobenius norm. For simplicity, one of the states is often chosen to be the $|0\rangle^{\otimes N}$ state. This process requires local (individual) qubit state tomography, which can be performed in parallel for each qubit to obtain the reduced density matrices, from which the most probable state is extracted.
@@
Concerning the optimization of parameters in the VQC, using the hyperspherical coordinates approach Bermejo et al. (2024), the cost function $f(\vec{x})$ and original $n$ Cartesian coordinates define a point $P$ in a $(n+1)$-dimensional space. The coordinates are then transformed into $(n+1)$-dimensional hyperspherical coordinates $\{\vec{\theta},r\}$ as follows:
@@ II.2 MPS Attack
    • BondDim: Virtual bond dimension of the MPS.
    • step_length: Learning rate for parameter updates.
    • steps: Number of sequential updates for the same tensor in each iteration.
    • cutoff_value: Threshold for discarding smaller eigenvalues in Singular Value Decomposition (SVD).
    • reset_value: Threshold for detecting local minima and resetting parameters.
    • temperature: Acceptance criterion for unfavorable changes.
@@
In this approach, the optimization is guided by the cost function, calculated as the Hamming Distance between the generated and target ciphertexts. The optimization process employs a simulated annealing method characterized by state and energy calculations, acceptance criteria based on the Metropolis-Hastings criterion, which allows the algorithm to escape local minima. The Adam optimizer is used to adapt the learning rate for each parameter, based on the first and second moments of the gradients, ensuring efficient convergence to the optimal solution.
@@ II.3 Flexible-PEPS Attack
The Flexible-PEPS based Quantum Circuit Simulator (FQCS) is a sophisticated simulation tool Patra et al. (2024b), leveraging TN methods to approximately simulate quantum circuits without an underlying lattice constraint. This section details the design, implementation, and advantages of FQCS in the context of quantum cryptographic analysis.
@@ III Results
In this section, we present the results of our cryptanalysis using tensor network methods on three symmetric-key ciphers. We benchmark these methods against traditional brute-force attacks and the VQAA to evaluate their efficiency and effectiveness in key recovery. We begin with S-DES, the most extensively studied cipher in our experiments, followed by S-AES and Blowfish.
@@ III.1 Simplified-Data Encryption Standard (S-DES)
A brute-force attack, requiring an average of 512 iterations, served as a baseline for comparison. In a brute-force attack, every possible key is tried until the correct one is found. Given that S-DES uses a 10-bit key, there are $10242^{10}=$ 1024 possible keys, and on average, half of these (512) would need to be tested to find the correct key. The brute-force approach took 2.51 seconds in our tests, which involved generating all possible keys using a Python script and encrypting the known plaintext with them. This time could be further reduced with dedicated hardware designed for such tasks.
@@ III.2 Simplified Advanced Encryption Standard (S-AES)
A brute-force attack, requiring an average of 32,768 iterations, served as a baseline for comparison. In a brute-force attack, every possible key is tried until the correct one is found. Given that S-AES uses a 16-bit key, there are $5362^{16}=65,$ 536 possible keys, and on average, half of these (32,768) would need to be tested to find the correct key. The brute-force approach took less than 14 seconds in our tests, which involved generating all possible keys using a Python script and encrypting the known plaintext with them. This time could be further reduced with dedicated hardware designed for such tasks.
@@
Again we found that using a bond dimension of 1 yielded faster results due to the algorithm’s simplicity in the MPS approach, but it was no longer more effective than FQCS in terms of the average number of iterations needed to recover the key. We believe the downgrade in performance could be due to over-parametrization (too many weights to optimize) if bond dimension is greater than 1, or lack of parameters in the case of bond dimension 1.
@@
Our analysis of S-AES cryptanalysis, summarized in Fig. 7, reveals that while MPS offers faster computation times compared to FQCS and VQAA due to its simplicity, it is not as efficient as brute-force methods in terms of speed. However, FQCS proves to be more effective in terms of the average number of iterations needed to recover the key, particularly when configured without CNOTs, effectively serving as a mean-field ansatz. This result underscores the potential of FQCS for tackling more complex ciphers with larger key sizes, highlighting the critical role of configuration optimization in achieving efficient quantum cryptanalysis.
@@ III.3 Blowfish
Given our previous experience with Blowfish in Ref.Aizpurua et al. (2023), instead of attempting a full attack on the 32-bit key, we employ a hybrid approach. We fix the first 8 bits of the key to the correct values, reducing the search space to the remaining 24 bits. This approach simulates running $2562^{8}=$ 256 configurations in parallel in few-qubit quantum processors. Our analysis used 6 qubits, implying the need for 256 independent 6-qubit quantum processors, which is realistic with current NISQ technology that handles thousands of qubits, such as those using neutral atoms Russell (2023). In this approach, a brute-force attack searching $2162^{24}=16,777,$ 216 possible keys would require on average half of these (8,388,608) to be tested to find the correct key.
@@ IV Conclusions
For S-DES, the results indicated that MPS with a bond dimension of 1 (product state) outperformed other methods in terms of iterations. The simplicity of S-DES and its small key size made it an ideal candidate for MPS, which efficiently updated tensors to recover the key. In the case of S-AES, the increase in key size and complexity led to longer runtimes. The MPS approach, while faster due to algorithm simplicity, was less effective than FQCS in terms of iterations needed to recover the key. For Blowfish, the hybrid approach of fixing the first 8 bits of the key reduced the search space to the remaining 24 bits, allowing for systematic analysis. The VQAA approach, using 6 qubits and incorporating CNOTs, was the most effective with an average of 4.3 million iterations and 15,558 seconds.
Overall, the study highlighted the potential of tensor network methods, particularly MPS and FQCS, in enhancing cryptographic key recovery. However, the effectiveness of these methods varied with the complexity of the cipher and the size of the key. VQAA demonstrated strong performance, particularly for more complex ciphers, emphasizing the importance of entanglement and the use of CNOTs in quantum cryptanalysis. While brute-force consistently achieved the fastest runtimes due to its simplicity, more advanced methods like FQCS and VQAA reduced the number of iterations required, especially for larger and more complex ciphers. As the complexity of the ciphers increased, there was a growing indication that a crossover point could occur where these more sophisticated methods would surpass brute force in runtime, making brute force no longer the fastest approach.
In a nutshell, the more complex the cipher and the longer the key, the more complex the best hacking method, as shown in Table 4.
@@
Acknowldgements.- We acknowledge Donostia International Physics Center (DIPC), Ikerbasque, Basque Government, Diputación de Gipuzkoa, European Innovation Council (EIC) and Tecnun for constant support, as well as insightful discussions with the teams from Multiverse Computing, DIPC and Tecnun on the algorithms and technical implementations. This work was supported by the Spanish Ministry of Science and Innovation through the project “Few-qubit quantum hardware, algorithms and codes, on photonic and solid-state systems” (PLEC2021-008251).

### v2

@@ I Introduction
Moreover, in the NISQ era, simulations can offer reliable benchmarks of noisy quantum computers Patra et al. (2024), which motivates us further to use our quantum-inspired tensor network methods. Such methods have also seen substantial advancements Orús (2014). Tensor networks such as Matrix Product States (MPS) have proven effective in simulating quantum systems Orús (2019); Verstraete et al. (2008). MPS are especially powerful in modeling one-dimensional systems, efficiently capturing their entanglement structure with polynomial resources. This efficiency stands in contrast to the exponential resources required by brute-force methods. The success of MPS in one-dimensional systems has inspired extensions to higher dimensions, such as Projected Entangled Pair States (PEPS), which handle two-dimensional systems and beyond. As an example, we recently introduced the so-called Flexible-PEPS approach Patra et al. (2025), able to simulate quantum complex systems without a geometry constraint. When simulating quantum computers, the resulting Flexible-PEPS Quantum Circuit Simulator (FQCS) is particularly powerful.
@@
In parallel with quantum and tensor-network-based cryptanalysis, recent cryptographic research has explored biologically inspired and hardware-driven encryption techniques, particularly for multimedia data. For example, chaotic systems and neural models have been employed in image and video encryption using 3D memristive cubic maps with dual discrete memristors Gao et al. (2025a), 2D Logistic-Rulkov neuron maps Gao et al. (2025b), and 2D extended Schaffer function maps combined with neural networks Gao et al. (2024). While these works target different applications—mainly focusing on data protection in visual and IoT contexts—they underscore the diversity of approaches emerging in modern cryptographic research and the growing importance of alternative computational models.
Recent quantum cryptanalysis efforts have explored both gate-based and annealing-based approaches. For instance, Phab et al. Phab et al. (2022) implemented the Quantum Approximate Optimization Algorithm (QAOA) on the Heys cipher using IBM’s NISQ devices, highlighting the limitations of current quantum circuit depth. Similarly, Pei et al. Pei et al. (2025) proposed a quantum annealing-based attack using D-Wave hardware on SPN ciphers, formulating the key-recovery problem as a constrained quadratic model. While these works demonstrate early hardware-based feasibility, they target toy ciphers and rely on quantum access. In contrast, our work introduces quantum-inspired tensor network methods—specifically, MPS and FQCS—which enable efficient classical simulation of quantum attacks on realistic symmetric-key protocols (S-DES, S-AES, Blowfish) without requiring access to quantum hardware. Our contributions lie in (i) algorithm design combining tensor networks with variational principles and Hamiltonian simulation, (ii) performance benchmarking across increasing cipher complexity, and (iii) analysis of scalability in entanglement, key size, and time/space cost.
@@ II.1 Variational Quantum Attack
where the columns of matrix $A$, i.e., $a_{i}$, represent the $k$ maximally orthogonal states over $N$ qubits and $\|\cdot\|_{F}$ denotes the Frobenius norm. For simplicity, one of the states is often chosen to be the $|0\rangle^{\otimes N}$ state. This process requires local (individual) qubit state tomography, which can be performed in parallel for each qubit to obtain the reduced density matrices, from which the most probable state is extracted. Note that in our work, these non-orthogonal states are not physically realized on quantum hardware, but rather classically emulated as part of our simulation framework.
@@
[…]
@@ II.2 MPS Attack
    • BondDim: Virtual bond dimension of the MPS (e.g., 20).
    • step_length: Learning rate for parameter updates (e.g., $10^{-2}$).
    • steps: Number of updates per tensor (e.g., 1).
    • cutoff_value: Threshold for SVD truncation (e.g., $10^{-8}$).
    • reset_value: Gradient norm threshold to reset parameters (e.g., 25).
    • temperature: Controls stochastic acceptance in updates (e.g., 1).
@@
In this approach, the optimization is guided by the cost function, calculated as the Hamming Distance between the generated and target ciphertexts. The optimization process employs a simulated annealing method characterized by state and energy calculations, acceptance criteria based on the Metropolis-Hastings criterion, which allows the algorithm to escape local minima. The Adam optimizer is used to adapt the learning rate for each parameter, based on the first and second moments of the gradients, ensuring efficient convergence to the optimal solution. We observed that moderate choices of bond dimension (1–25) and step sizes in the range $10^{-3}$$10^{-2}$ generally achieved good performance across all tested ciphers. The optimization was robust to small variations in hyperparameters, and excessive tuning was not required.
@@ II.3 Flexible-PEPS Attack
The Flexible-PEPS based Quantum Circuit Simulator (FQCS) is a sophisticated simulation tool Patra et al. (2025), leveraging TN methods to approximately simulate quantum circuits without an underlying lattice constraint. This section details the design, implementation, and advantages of FQCS in the context of quantum cryptographic analysis.
@@
A detailed step-by-step flowchart of the FQCS simulation algorithm is provided in Appendix A, outlining the sequence of gate application, entanglement update, and truncation procedures.
@@
The main tunable parameters for FQCS include the same as in the MPS attack, adding number_layers (typically 2) and number_qubits (e.g., 16) for the circuit simulation. As with the MPS attack, performance was robust across reasonable values, with convergence most sensitive to the number of layers and the reset threshold. Unlike in MPS, we did not vary the physical bond dimension extensively in FQCS, as it mostly affects the precision of tensor contractions. With relatively shallow circuits and few qubits, we observed minimal performance differences across bond dimension settings.
@@
Although FQCS is designed as a classical simulation framework, the quantum circuits it simulates consist entirely of standard one- and two-qubit gates, which are natively supported by current quantum hardware. This makes FQCS-generated circuits directly implementable on gate-based quantum platforms. However, the depth and entanglement of the circuits must be carefully managed to remain within the limits imposed by decoherence and gate noise. In this context, FQCS serves as a valuable tool for prototyping and benchmarking variational quantum attacks prior to deployment on real devices.
@@ III Results
In this section, we present the results of our cryptanalysis using tensor network methods on three symmetric-key ciphers. We benchmark these methods against traditional brute-force attacks and the VQAA to evaluate their efficiency and effectiveness in key recovery. We begin with S-DES, the most extensively studied cipher in our experiments, followed by S-AES and Blowfish. To ensure statistical significance, all reported averages are computed over a large number of independent attack attempts (e.g., 200 for S-DES and 100 for S-AES and Blowfish), each with randomized key and plaintext-ciphertext pairs.
@@ III.1 Simplified-Data Encryption Standard (S-DES)
A brute-force attack, requiring an average of 512 iterations, served as a baseline for comparison. In a brute-force attack, every possible key is tried until the correct one is found. Given that S-DES uses a 10-bit key, there are $2^{10}=1024$ possible keys, and on average, half of these (512) would need to be tested to find the correct key. The brute-force approach took 2.51 seconds in our tests, which involved generating all possible keys using a Python script and encrypting the known plaintext with them. This time could be further reduced with dedicated hardware designed for such tasks.
@@ III.2 Simplified Advanced Encryption Standard (S-AES)
A brute-force attack, requiring an average of 32,768 iterations, served as a baseline for comparison. In a brute-force attack, every possible key is tried until the correct one is found. Given that S-AES uses a 16-bit key, there are $2^{16}=65,536$ possible keys, and on average, half of these (32,768) would need to be tested to find the correct key. The brute-force approach took less than 14 seconds in our tests, which involved generating all possible keys using a Python script and encrypting the known plaintext with them. This time could be further reduced with dedicated hardware designed for such tasks.
@@
Again we found that using a bond dimension of 1 yielded faster results due to the algorithm’s simplicity in the MPS approach, but it was no longer more effective than FQCS in terms of the average number of iterations needed to recover the key. We believe the downgrade in performance could be due to over-parametrization (too many weights to optimize) if bond dimension is greater than 1, or lack of parameters in the case of bond dimension 1. We also tested higher bond dimensions and observed that while they generally reduced the number of iterations required, they increased runtime due to the added computational overhead. As such, a bond dimension of 1 provided the best trade-off between accuracy and speed for small key sizes.
@@
Our analysis of S-AES cryptanalysis, summarized in Fig. 7, reveals that while MPS offers faster computation times compared to FQCS and VQAA due to its algorithmic simplicity, it is not as efficient as brute-force methods in terms of raw speed. Among quantum-inspired approaches, FQCS proves more effective in terms of the average number of iterations needed to recover the key, particularly when configured without CNOTs—effectively serving as a mean-field ansatz. This highlights the importance of optimizing circuit configurations for practical quantum cryptanalysis.
A direct comparison between MPS and VQAA-h further clarifies this trade-off. As shown in Fig. 7, VQAA-h outperforms MPS in terms of average iterations required to recover the key. However, MPS remains significantly faster overall due to its lower per-iteration cost and simpler tensor updates. This reflects a core distinction: VQAA-h, simulated via FQCS, accesses a richer variational space incorporating entanglement, but at higher computational expense. In contrast, MPS—especially at low bond dimension—offers rapid convergence with reduced expressivity. As cipher complexity and key size increase, we anticipate that the advantages of VQAA-h may become more pronounced, provided sufficient simulation resources are available.
@@ III.3 Blowfish
Given our previous experience with Blowfish in Ref. Aizpurua et al. (2023), instead of attempting a full attack on the 32-bit key, we employ a hybrid approach. We fix the first 8 bits of the key to the correct values, reducing the search space to the remaining 24 bits. This approach simulates running $2^{8}=256$ configurations in parallel in few-qubit quantum processors. Our analysis used 6 qubits, implying the need for 256 independent 6-qubit quantum processors, which is realistic with current NISQ technology that handles thousands of qubits, such as those using neutral atoms Russell (2023). In this approach, a brute-force attack searching $2^{24}=16,777,216$ possible keys would require on average half of these (8,388,608) to be tested to find the correct key.
@@
For the MPS approach, the per-sweep time complexity is $\mathcal{O}(D^{3})$ and the space complexity is $\mathcal{O}(D^{2})$, where $D$ is the bond dimension. While small bond dimensions (e.g., $D=1$) are sufficient for simple ciphers like S-DES, larger values of $D$ are needed to capture more entanglement in complex ciphers, which significantly increases runtime and memory requirements.
The time complexity of the FQCS algorithm is $\mathcal{O}(D^{2\kappa-1})$, and space complexity $\mathcal{O}(D^{\kappa})$, where $\kappa$ is the max vertex degree. Although FQCS benefits from entanglement pruning, high $\kappa$ values and deep circuits for 64- or 128-bit keys demand large memory and compute resources. In practice, VQAA-h offers better key recovery with fewer iterations for intermediate-size ciphers (e.g., S-AES), but the higher computational overhead makes MPS more practical for quick attacks on simpler protocols (e.g., S-DES).
VQAA simulations, based on full statevector backends, have exponential time and space complexity $\mathcal{O}(2^{N})$, limiting practical runs to fewer than 25 qubits. In contrast, brute-force attacks scale as $\mathcal{O}(2^{k})$ in time and remain infeasible beyond 64-bit keys.
Overall, while all methods face scaling challenges, tensor network approaches (MPS, FQCS) offer more tractable classical alternatives for intermediate key lengths. Further optimization or parallelization would be required for full 64/128-bit cryptanalysis.
@@ IV Conclusions
For S-DES, the results indicated that MPS with a bond dimension of 1 (product state) performed remarkably well, benefiting from the simplicity of the cipher and the limited entanglement required. This low-complexity setup allowed MPS to achieve high efficiency in both convergence and runtime. In contrast, for the more intricate S-AES cipher, FQCS showed stronger performance in terms of average iterations, despite longer runtimes. The ability of FQCS to approximate more expressive circuit structures—while remaining classically executable—proved beneficial in this intermediate regime. Finally, for Blowfish, the VQAA method, incorporating entangling CNOT gates and simulated over six qubits, achieved the best key recovery rates. This reflects the increasing relevance of quantum resources as cipher complexity scales.
Overall, these findings suggest that the balance between algorithmic expressivity and computational cost plays a critical role in cryptanalysis. While brute-force attacks remain effective for small key spaces, more sophisticated methods such as MPS, FQCS, and VQAA offer advantages in scalability and search space exploration. The progression of results across the three ciphers hints at a gradual transition: as key size and internal cipher structure become more complex, the effectiveness of quantum-inspired and tensor-based methods becomes increasingly evident.
This evolution is summarized in Table 4, which aligns each cipher with the attack method that yielded the most favorable trade-off between efficiency and complexity.
@@
Acknowldgements.- We acknowledge Donostia International Physics Center (DIPC), Ikerbasque, Basque Government, Diputación de Gipuzkoa, European Innovation Council (EIC) and Tecnun for constant support, as well as insightful discussions with the teams from Multiverse Computing, DIPC and Tecnun on the algorithms and technical implementations. This work was supported by the Spanish Ministry of Science and Innovation through the project “Few-qubit quantum hardware, algorithms and codes, on photonic and solid-state systems” (PLEC2021-008251) and by the Diputación Foral de Gipuzkoa through the “Biased quantum error mitigation and applications of quantum computing to gene regulatory networks” project (2024-QUAN-000020).
Data availability.- Te datasets used and/or analysed during the current study are available from the corresponding author on reasonable request.
Appendix A Flowchart for FQCS Simulation Algorithm
In this appendix, we present a detailed flowchart of our Flexible PEPS-based Quantum Circuit Simulator (FQCS). The primary goal of FQCS is to simulate quantum circuits composed of arbitrary one- and two-body gates. The simulation backend is based on the Flexible Projected Entangled Pair States (fPEPS) tensor network algorithm Patra et al. [2025].
The fPEPS framework captures entanglement between qubits via virtual bonds connecting vertex tensors that represent the individual qubit states. A higher virtual bond dimension corresponds to higher entanglement content. To maintain computational efficiency—even for non-local gates and densely connected circuits—the method dynamically prunes virtual edges based on the instantaneous correlation structure (quantified by bond entanglement entropy). FQCS leverages this fPEPS formalism to apply gate sequences to a tensor network representing the evolving quantum state. Below, we outline the core steps of the FQCS simulation in a structured flowchart format, encompassing both gate application and entanglement management.
High-Level Flowchart of FQCS (with Hyperparameters $\chi$ and $\kappa$)
  • Initialization: Initialize a product state on $N$ qubits/spins by assigning a vertex tensor $\Gamma_{i}$ (a vector) to each site $i$. The initial tensor network contains no virtual bonds. This corresponds to the quantum circuit initialization where all qubits are in the state $|0\rangle_{i}$.
  • For each gate $O$ in the quantum circuit:
    • If $O$ is a one-body gate $O_{i}$:
      Apply the gate via exact local matrix multiplication:
      $\Gamma_{i}\leftarrow O_{i}\cdot\Gamma_{i}$
    • If $O$ is a two-body gate $O_{ij}$:
      1. 1.
        Apply the gate using the Simple Update method: apply $O_{ij}$ locally to the vertex tensors $\Gamma_{i}$ and $\Gamma_{j}$, followed by singular value decomposition to obtain updated vertex tensors and an edge tensor $\lambda_{ij}$ (or update the existing edge tensor if one already exists).
      2. 2.
        Truncate the bond dimension of $\lambda_{ij}$ to a maximum value $\chi$, preserving only the dominant singular values.
      3. 3.
        Update $\Gamma_{i}$, $\Gamma_{j}$, and $\lambda_{ij}$ accordingly after truncation.
      4. 4.
        Check: Does either vertex ($i$ or $j$) now have more than $\kappa$ virtual edges?
      5. 5.
        If yes: For each overconnected vertex $x\in\{i,j\}$
        1. (a)
[…]
        2. (b)
          Sort the BEE values to obtain a ranked list ¯$\bar{\mathcal{E}}_{x}$.
        3. (c)
          Identify the edge $(x,y)$ with the lowest BEE to truncate.
        4. (d)
          Truncate edge $(x,y)$ as follows:
            Reshape the vertex tensors $\Gamma_{x}$ and $\Gamma_{y}$ into matrices $M_{x}$ and $M_{y}$.
            Perform a rank-1 truncation (matrix slicing) by retaining only the leading singular value in $\lambda_{xy}$.
            Absorb $\sqrt{\lambda_{xy}}$ into $M_{x}$ and $M_{y}$:
            $M_{x}^{\prime\prime}=M_{x}^{\prime}\cdot\sqrt{\lambda_{xy}},\quad M_{y}^{\prime\prime}=\sqrt{\lambda_{xy}}\cdot M_{y}^{\prime}$
            Reshape $M_{x}^{\prime\prime}$ and $M_{y}^{\prime\prime}$ back into updated vertex tensors $\Gamma_{x}^{\prime}$ and $\Gamma_{y}^{\prime}$.
            Remove the edge $(x,y)$ from the network.
  • Repeat: Iterate over all gates in the circuit.
The final tensor network represents an approximate simulation of the full quantum circuit, governed by the approximation parameters $\chi$ and $\kappa$, where $\chi,\kappa\in\mathbb{Z}^{+}$. Larger values of these parameters allow more accurate simulations at the cost of increased computational complexity.